Cimetidine is used in the treatment and management of gastric ulcers and gastroesophageal reflux disease. In the case of an active ulcer, cimetidine can be administered continuous by IV infusion at a dose rate of 37.5 mg/hr. If the concentration of the cimetidine solution is 6 mg/ml, then how many milliliters must be available for a 4-hr infusion?

Misoprostol is used to prevent NSAID-induced gastric ulcers. A daily dose consists of two tablets administered four times each day. What is the size of one day’s dose if each tablet contains 100 mcg of misoprostol?

Omeprazole is widely used to treat gastroesophageal reflux disease (GERD). A daily dose can consist of a single dose of 40 mg/day. How many days of dosage will a bottle of 48 tablets, each containing 20 mg of omeprazole, provide?

Used to treat different forms of nausea, promethazine is typically administered every 6 hours at individual doses of 12.5 mg. How much promethazine is there in a week’s dose?
Answer #1

ANSWER.

* Calculation of how many ml of cimetidine available for 4hr infusion.

            Formula = Desired dose / Dose in hand * Quantity.

              Desired dose (D) = 37.5 mg /hr

              Dose in hand (H) = 6 mg.

                         Quantity(Q) = 1ml.

        As per formula

             Amount(ml) of drug available = D/H *Q

                                                            = 37.5/6 * 1

                                                            = 6.25 * 1

                                                            = 6.25 ml/hr.

           Amount (ml) of drug required for 4 hr = 6.25 * 4

                                                                       = 25 ml.

* Size of the one day dose of misoprostol.

            Dose ordered per day = 2 tablets 4 times each day.(q6h)

               Size of one tablet of misoprostol = 100 mcg.

            Number of tablets required for one day = 4 * 2

                                                                           = 8 tablets.

            Size of the tablets required for one day = 8 * 100

                                                                          = 800 mcg /day.

* Calculation of how many days of dosage with a bottle of 48 tablets.

            Ordered dose of omeprazole = 40 mg / day.

            Number of tablets in the bottle = 48 tablets.

           Dose of each tablet in the bottle = 20 mg.

      Total amount of dosage provide by 48 tablets = 48 * 20

                                                                               = 960 mg.

       Number of days of dosage provide by 48 tablets = 960 /40

                                                                                     = 24 days.

* Amount of promethazine in a week's dose.

      Ordered dose of promethazine = 12.5 mg every 6 hr.(4 times a day)

        Dose of promethazine required for one day(24 hr) = 12.5 * 4

                                                                                         = 50 mg.

         Amount of promethazine in a week's dose(7 days) = 50 * 7

                                                                                 = 350 mg.